# Detrimental non-Markovian errors for surface code memory

### v2

@@ IV.2 Varying correlated noise model parameters
Figure 6: Teraquop distance projection plots for polynomially and exponentially decaying streaky error models for Class 1 and Class 2 only errors. We test different values of $n$ in $Ap/t^{n}$ and $Ap/n^{t}$ for the polynomial and exponential model, respectively. Models with $n\to\infty$ are essentially restricted to streak lengths of only two rounds for polynomial decay models or no streaks for exponentially decay models.
@@ V.3 Conclusion
We have explored the impact of temporally correlated error structures, based off circuit-level noise, on the performance of surface code memory. Our analysis has found that while fault tolerance is realistically achievable under arbitrary long-range two-time correlated errors, multi-time “streaky” correlated errors involving syndrome qubits can significantly degrade logical error rate scaling. Specifically, we found that quadratically decaying streaky correlations leads to power law scaling of logical error rates, with no feasible teraquop distance projections.
@@ Appendix A Error marginalization
In this section, we outline our method for obtaining marginal error probabilities from a temporally correlated noise model.
Define $\mathcal{T}$ to be the set of all possible error events $E_{i}$ that each occur with some independent probability $P_{i}$:
$\mathcal{T}=\{E_{1},E_{2},\dots,E_{n}\},$ (6)
$\mathrm{Pr}(E_{i})=P_{i}.$ (7)
Each error event $E_{i}$ is a set of mutually exclusive error configurations $E_{i}=\{\mathbf{X}^{(i)}_{1},\mathbf{X}^{(i)}_{2},\cdots,\mathbf{X}^{(i)}_{m}\}$ associated with a probability distribution $\mathrm{Pr}(\mathbf{X}^{(i)}_{j}|E_{i})$, where $\sum_{j}\mathrm{Pr}(\mathbf{X}^{(i)}_{j}|E_{i})=1$. Error configurations $\mathbf{X}^{(i)}_{j}$ represent sites where an error has or hasn’t occurred. For our purposes, since we only deal with single Pauli-flip or uniform depolarizing error channels, $\mathbf{X}^{(i)}_{j}$ corresponds to a binary string, where 1 denotes an error, and 0 denotes no error.
Error marginals $p(\mathcal{E},\mathbf{x})$ define the strength of the noise channel after each gate in a circuit-level noise model, where $\mathcal{E}$ denotes one of the channels defined in Eq. 1, and $\mathbf{x}$ the location of the error in the circuit. To calculate $p(\mathcal{E},\mathbf{x})$, we first calculate the marginal probabilities $q_{i}(\mathcal{E},\mathbf{x})$ due to each individual error event $E_{i}$:
$q_{i}(\mathcal{E},\mathbf{x})=P_{i}\sum_{j|\mathbf{X}^{(i)}_{j}(\mathbf{x})\neq I}\mathrm{Pr}(\mathbf{X}^{(i)}_{j}|\mathcal{E}_{i}).$ (8)
To combine each of the $q_{i}(\mathcal{E},\mathbf{x})$ probabilities to obtain the final marginal probability, while accounting for cancellation of errors, we perform the product
$1-t(\mathcal{E},\mathbf{x})=\prod_{i}\left(1-t_{i}(\mathcal{E},\mathbf{x})\right),$ (9)
where
$t(\mathcal{E}^{1}_{X},\mathbf{x})=2p(\mathcal{E}^{1}_{X},\mathbf{x})$ (10)
$t_{i}(\mathcal{E}^{1}_{X},\mathbf{x})=2q_{i}(\mathcal{E}^{1}_{X},\mathbf{x})$ (11)
$t(\mathcal{E}^{1}_{D},\mathbf{x})=\frac{4}{3}p(\mathcal{E}^{1}_{D},\mathbf{x})$ (12)
$t_{i}(\mathcal{E}^{1}_{D},\mathbf{x})=\frac{4}{3}q_{i}(\mathcal{E}^{1}_{D},\mathbf{x})$ (13)
$t(\mathcal{E}^{2}_{D},\mathbf{x})=\frac{16}{15}p(\mathcal{E}^{2}_{D},\mathbf{x})$ (14)
$t_{i}(\mathcal{E}^{2}_{D},\mathbf{x})=\frac{16}{15}q_{i}(\mathcal{E}^{2}_{D},\mathbf{x}).$ (15)
Specifically, $p(\mathcal{E}^{1}_{X},\mathbf{x})$ is obtained using Eqs. 9, 10 and 11, $p(\mathcal{E}^{1}_{D},\mathbf{x})$ using Eqs. 9, 12 and 13, and $p(\mathcal{E}^{2}_{D},\mathbf{x})$ using Eqs. 9, 14 and 15.
@@
Error Class Structure Decay Function Parameters($A,p,n$) Fit Teraquop Distance
@@
Polynomial: $Ap/\Delta t^{n}$, Exponential: $Ap/n^{\Delta t}$, *No realistic teraquop projection

### v3

@@ IV.2 Varying correlated noise model parameters
Figure 6: Teraquop distance projection plots for polynomially and exponentially decaying streaky error models for Class 1 and Class 2 only errors. We test different values of $n$ in $Aq/t^{n}$ and $Aq/n^{t}$ for the polynomial and exponential model, respectively. Models with $n\to\infty$ are essentially restricted to streak lengths of only two rounds for polynomial decay models or no streaks for exponentially decay models.
@@ V.3 Conclusion
We have explored the impact of temporally correlated error structures, based off circuit-level noise, on the performance of surface code memory. Our analysis has found that while fault tolerance is realistically achievable under arbitrary long-range two-time correlated errors, multi-time “streaky” correlated errors involving syndrome qubits can significantly degrade logical error rate scaling. Specifically, extrapolating simulation data on logical error rates, we find that quadratically decaying streaky correlations result in, at best, slower-than-exponential suppression, with no feasible distance projection for achieving teraquop error rates.
@@ Appendix A Error marginalization
In this appendix, we outline our method for obtaining marginal error probabilities from a temporally correlated noise model. We then provide a concrete, worked example of the marginalization process for a pairwise time-correlated correlated error model.
For convenience, we begin by reparametrizing the error channels given in Eq. 1 as follows.
$\mathcal{E}^{1}_{X}(p^{\prime}{,\rho})=(1-p^{\prime})\rho+\frac{p^{\prime}}{2}\left(\rho+X\rho X\right)$ (6a)
$\mathcal{E}^{1}_{D}(p^{\prime}{,\rho})=(1-p^{\prime})\rho+\frac{p^{\prime}}{4}(\rho+X\rho X+Y\rho Y+Z\rho Z)$ (6b)
$\mathcal{E}^{2}_{D}(p^{\prime}{,\rho})=(1-p^{\prime})\rho+\frac{p^{\prime}}{16}\sum_{P\in\{I,X,Y,Z\}^{\otimes 2}}P\rho P.$ (6c)
Each of these channels is equivalent to the corresponding channel in Eq. 1 under the transformation $p^{\prime}=C_{\mathcal{E}}p$, where $C_{\mathcal{E}}$ is a constant that depends on the error channel $\mathcal{E}$. Specifically,
$C_{\mathcal{E}}=\begin{cases}2,&\mathcal{E}=\mathcal{E}^{1}_{X}\\ \frac{4}{3},&\mathcal{E}=\mathcal{E}^{1}_{D}\\ \frac{16}{15},&\mathcal{E}=\mathcal{E}^{2}_{D}\\ \end{cases}.$ (7)
This moves away from the standard Stim uses (Eq. 1) to the equivalent interpretation in which a channel $\mathcal{E}(p^{\prime},\rho)$ maximally “mixes” a state $\rho$ with probability $p^{\prime}$, where the maximally mixing channel $\mathcal{M}$ has the convenient property $(\mathcal{M}\circ\mathcal{M})[\rho]=\mathcal{M}[\rho]$. Consequently, we define $p^{\prime}(s,t)$ as the probability that qubit $s$ at time $t$ is maximally mixed by the noise channel. To compute the probability $p(s,t)$, we first compute $p^{\prime}(s,t)$ and then apply the conversion using the factor $C_{\mathcal{E}}$, where $\mathcal{E}$ implicitly depends on both $s$ and $t$ (i.e., what gate immediately precedes the error channel).
We reparametrize the channels in this way to enable the use of the following formula for calculating the probability that an error occurs at any given space-time coordinate in the experiment. If qubit $s$ at time $t$ can experience $k$ possible errors, each of which maximally mixes the qubit and occurs independently with probability $p_{i}^{\prime}$, for $i\in{1,\dots,k}$, then the probability that no mixing occurs is given by:
$1-p^{\prime}(s,t)=\prod_{i=1}^{k}(1-p_{i}^{\prime}).$ (8)
This formula is based on the property that applying two maximally mixing channels has the same effect as applying a single maximally mixing channel: $(\mathcal{M}\circ\mathcal{M})[\rho]=\mathcal{M}[\rho]$. Thus, the probability that the qubit does not become mixed is simply the probability that none of the $k$ “mixing” errors occur. This holds for single Pauli flip errors as well, where $p^{\prime}(s,t)$ is twice the probability that qubit $s$ at time $t$ experiences a Pauli flip.
Define $\mathcal{T}$ as the set of all error events that can occur during the experiment. Each event $E\in\mathcal{T}$ occurs independently with probability $\mathrm{Pr}(E)$. This event does not necessarily correspond to a single error channel but rather to a set of possible error channels $\{X_{1}^{E},\dots,X^{E}_{m}\}$, with one channel selected with probability $\mathbb{P}\left(X_{i}^{E}\mid E\right)$.
For the models considered in this paper, the set of all error events affecting qubit $s$ is denoted as $\mathcal{T}_{s}:=\{E_{s,i,j}~{}|~{}1\leq i<j\leq N\}$, where $N$ is the number of rounds in the experiment; it follows that $\mathcal{T}$ is simply the (disjoint) union of $\mathcal{T}_{s}$ for all $s$. In the pairwise model, the set of errors corresponding to $E_{s,i,j}$ is the set of all Pauli strings on qubit $s$ at times $i$ and $j$, i.e.,
$\{I_{i},X_{i},Y_{i},Z_{i}\}\otimes\{I_{j},X_{j},Y_{j},Z_{j}\}.$ (9)
In the streaky model, the set of errors is given by
$\bigotimes_{k=i}^{j}\{I_{k},X_{k},Y_{k},Z_{k}\}.$ (10)
In both models, the corresponding distribution $\mathbb{P}$ over these sets is simply the uniform distribution. That is, if error event $E_{s,i,j}$ occurs, qubit $s$ experiences a uniformly random Pauli error (including the possibility of the identity) at times $i$ and $j$ (and also times $i\leq t\leq j$ in the case of the streaky model). Then, for the pairwise model, it follows from Eq. 8 that $p^{\prime}(s,t)$, the probability that qubit $s$ is maximally mixed at time $t$, satisfies
$1-p^{\prime}(s,t)=\prod_{1\leq i<j\leq N}\left(1-\mathrm{Pr}\left(E_{s,i,j}\right)\mathbbm{1}\{t\in\{i,j\}\}\right),$ (11)
where $\mathbbm{1}\{t\in\{i,j\}\}$ is $1$ if $t\in\{i,j\}$ and 00 otherwise. In the streaky model, we similarly obtain
$1-p^{\prime}(s,t)=\prod_{1\leq i<j\leq N}\left(1-\mathrm{Pr}\left(E_{s,i,j}\right)\mathbbm{1}\{t\in[i,j]\}\right).$ (12)
In the example below we write $u_{ij}(t)=\mathrm{Pr}\left(E_{s,i,j}\right)\mathbbm{1}\{t\in\{i,j\}\}$ for a fixed $s$. Thus, $p(s,t)$ is computed by
$p(s,t)=\frac{1}{C_{\mathcal{E}}}\left[1-\prod_{1\leq i<j\leq N}\left(1-u_{ij}(t)\right)\right].$ (13)
Example: Pairwise correlated depolarizing errors
Consider a pairwise time-correlated depolarizing error model on a single qubit $s$ over $N$ timesteps. We wish to write the marginalized probability distribution $p(s,t)=p(t)$ for $t\in\{1,2,\dots,N\}$ (suppressing $s$ in the notation as we focus on just a single qubit), where $p(t)$ is the probability the qubit experiencing a uniformly random non-identity Pauli error at timestep $t$ — i.e., $\mathcal{E}(t)=\mathcal{E}^{1}_{D}(p(t))$. Define the probability that 1 of 15 equally likely pairs of errors $IX,IY,IZ,\dots,ZZ$ occur at timesteps $i$ and $j$ to be $Aq/|i-j|^{n}$. This matches the definition of the pairwise error model given in Section III.1. Then the probability the qubit at timestep $t$ becomes maximally mixed (equal chance $I,X,Y,Z$) due to each error event $E_{ij}|t\in\{i,j\}$ is
$u_{ij}(t)=\frac{16}{15}\frac{Aq}{|i-j|^{n}}.$ (14)
The factor of $16/15$ is due to how we defined our error model in the main text, where we followed the convention Stim uses to define Pauli error channels. Thus, inputting Eq. 14 into Eq. 13 and taking $C_{\mathcal{E}}=\frac{4}{3}$, we obtain
$p(t)=\frac{3}{4}\left[1-\prod_{1\leq i<j\leq N}\left(1-\frac{16Aq}{15|i-j|^{n}}\right)\right].$ (15)
Appendix B Non-monotonicity of logical error rate scaling under detrimental noise models
In the main text, our analysis of the streaky correlated noise model fitted the logical error rate to a power-law decay, as shown in the bottom panel of Fig. 3. However, this fit does not preclude the possibility of more pathological behavior. Small but systematic deviations between the data and the power-law trend suggest that the observed scaling may not persist at larger code distances. In particular, the logical error rate could exhibit non-monotonic behavior—initially decreasing with code distance, but eventually saturating or even increasing at larger distances.
To further probe this behavior, we performed additional simulations of the streaky correlated model while varying the characteristic correlated error probability $q$. The results, shown in Fig. B1, reveal instances of non-monotonic logical error rate scaling. For $q=2\times 10^{-3}$, the logical error rate at $d=11$ exceeds that at $d=13$; similarly, for $q=3\times 10^{-3}$, the logical error rate at $d=11$ exceeds that at $d=9$, further increasing with code distance. It is worth noting that for $q=3\times 10^{-3}$, the marginal physical error rate approaches the surface code threshold, which also leads to degraded performance in the corresponding marginalized independent model—though this degradation remains monotonic.
These results provide evidence that, for sufficiently high physical error rates and code distances, the logical error rate under streaky correlated noise can exhibit non-monotonic behavior—even when the corresponding marginalized independent noise model continues to show the expected monotonic suppression. Whether such non-monotonic scaling persists across all values of the characteristic correlated parameter $q$ remains an open question and may require rigorous analytical treatment to resolve. Regarding the existence of a threshold, we emphasize that even under monotonic decay, the absence of exponential suppression is sufficient to disqualify a true threshold as conventionally defined [38]. This is why we refer to an “apparent threshold” in the main text. Altogether, these findings reinforce our central conclusion: certain temporally correlated noise structures can fundamentally compromise the fault-tolerance properties of surface codes when using a standard decoder.
Figure B1: Logical error rate per round versus surface code distance under the streaky correlated model. Data is shown for several correlated error parameters. The data for $q=1\times 10^{-3}$ corresponds to the data shown in the bottom panel of Fig. 3. For higher values of $q$, the logical error rate scaling is non-monotonic, reaching a minimum before increasing at larger code distances. We also display data for the corresponding marginalized error models for parameters $q=0.002$ and $q=0.003$. Of note is the marginalized model data for $q=0.002$ which indicates exponential suppression of logical error rate for uncorrelated errors in contrast to the non-monotonic behavior of the corresponding correlated model.
@@ Appendix B Non-monotonicity of logical error rate scaling under detrimental noise models
Error Class Structure Decay Function Parameters($A,q,n$) Fit Teraquop Distance
@@
Polynomial: $Aq/\Delta t^{n}$, Exponential: $Aq/n^{\Delta t}$, *No realistic teraquop projection